Jan is choosing between three alternative investments, each of which will require a $100,000 initial outlay. Determine the present value of the after-tax cash flows from investment B described below.

Assume the following in your computations:

  • Jan is in the 24% marginal bracket; Long-Term Capital Gains rate is 15%;
  • All tax payments occur at the end of the year;
  • 4% discount rate (present value interest factors are below).

N

PV $1

PV Annuity

1

0.962

0.962

2

0.925

1.887

3

0.889

2.776

Investment B is a corporate bond that will pay 7% interest at the end of each year for 3 years, with the principal recovered at the end of the 3rd year. (the answer is $3,668, but can you show the work)
Answer #1

We will find after tax interest cash inflow

100000*7%=7000$

Tax =7000*24%

=1680$ per year

After tax cash inflow=7000-1680=

Years cash flow Pv factor present value
0 (100000) 1 (100000)
1-3 5320 2.776 (use annuity as the interest income is for three year) 14768(5320*2.776)
3 100000 0.889 88900
Net present value 3668$(14768+88900-100000)